OpenStudy (anonymous):

i solved 7x^2 – 63 = 0 which gave me x=9. How do i find the other x?

OpenStudy (wasiqss):

u did it wrong actually x^2=9, it means 2 values must satisfy this equation, these values are +3 and -3

OpenStudy (anonymous):

Yeah, like wasiqss said. There's always two possible solutions for a squared value, since squaring a negative returns the same result as squaring the same positive - i.e. 3^2 = (-3)^2 = 9.

OpenStudy (anonymous):

Well, "always" if a value isn't 0.

OpenStudy (anonymous):

so i should have squared it after i got the 9?

OpenStudy (anonymous):

why is it two 3's instead of one?

OpenStudy (dumbcow):

whenever you take the square root, there is a plus-minus condition x^2 = n --> x = +-sqrt(n) This is because the square of a negative number is positive x^2 = 9 --> x = +-3 3^2 = 9 (-3)^2 = 9

OpenStudy (anonymous):

thanks!